OPTICAL-OPTICAL DOUBLE RESONANCE SPECTROSCOPY OF YTTRIUM MONOHALIDES

\underline{JIANJUN YE}, H.F. PANG, AND A.S-C. CHEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.

High resolution optical-optical double resonance (OODR) spectra of YCl and YBr in the spectral region between 338.9 and 371.7 nm have been obtained using laser radiation from continuous wave dye and Ti:sapphire lasers pumped by argon ion lasers. The OODR spectrum was observed by recording the laser induced fluorescence from the excited state. Reacting laser-ablated yttrium atoms, respectively, with BCl_{3} and C_{2}H_{5}Br seeded in helium produced YCl and YBr molecules. For YCl, the [27.2] ^{1}Δ state was reached via the intermediate B^{1}Π from the X^{1}Σ state. The molecular constants for the v = 1 level of the [27.2] ^{1}Δ state were determined. For YBr, two new electronic states, namely: [26.0] ^{1}Π and [29.0] ^{1}Π were observed via the intermediate C^{1}Σ state from the X^{1}Σ state. Accurate molecular constants for the v = 4 and 5 levels of the [26.0] ^{1}Π state and the v = 1 and 2 of the [29.0] ^{1}Π state were determined. The observation of the spectra of isotopic molecules confirmed the vibrational quantum number assignment of the measured vibronic levels.